OpenStudy (anonymous):

What is the polynomial function of lowest degree with lead coefficient 1 and roots i, –2, and 2?

OpenStudy (dumbcow):

best approach is to set "x" equal to each root. x = i x = -2 x = 2 then set each one equal to 0, always square the equations containing "i" x^2 = i^2 = -1 ----> x^2 + 1 = 0 x+2 = 0 x-2 = 0 Finally put all the factors together... ---> (x^2+1)(x+2)(x-2)